• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name C. C. Chang

    Log entry time 00:02:12 on June18,2002

    Entry number 86314

    keyword=Shift Summary

    16:15 Q1 crashed.
    16:32 While current on Q1 is being ramped up passed 2000 A (but below
    desired set current of 2695 A), it tripped again. Called Scott Spiegel
    and he is on his way to make an access to the hall to investigate.
    17:17 Scott claimed that the He level is too low (at about 66%) and at 2695 A
    we might have He boils off causing He level to go below trip level,
    thus causing the trip. He requests to increase the He cooling level to 70%.
    We manage to reset Q1 current and resume data taking.
    20:15 At the end of Run 1599 we had 663 coincidence events in 358 muA hr.
    The run plan calls at least 800 events. After talking to Andi, the RC,
    we decided to have another run at Kinematics Q2_f30r which should
    give us additional 160 - 200 events. After taking the efficiency run
    we will switch to Kinematics Q2_f30l which should occur at around
    20:30.
    21:05 Communication between CODA and ROCs broken. Just Reset and
    Download CODA to restore the connection.
    24:00 In the process of taking Kinematics Q2_f30l. It has been a good shift.